# Audio Visual Segmentation Through Text Embeddings

### v1

@@ Abstract
The goal of Audio-Visual Segmentation (AVS) is to localize and segment the sounding source objects from the video frames. Researchers working on AVS suffer from limited datasets because hand-crafted annotation is expensive. Recent works attempt to overcome the challenge of limited data by leveraging the segmentation foundation model, SAM, prompting it with audio to enhance its ability to segment sounding source objects. While this approach alleviates the model’s burden on understanding visual modality by utilizing pre-trained knowledge of SAM, it does not address the fundamental challenge of the limited dataset for learning audio-visual relationships. To address these limitations, we propose AV2T-SAM, a novel framework that bridges audio features with the text embedding space of pre-trained text-prompted SAM. Our method leverages multimodal correspondence learned from rich text-image paired datasets to enhance audio-visual alignment. Furthermore, we introduce a novel feature, , which emphasizes shared semantics of audio and visual modalities while filtering irrelevant noise. Experiments on the AVSBench dataset demonstrate state-of-the-art performance on both datasets of AVSBench. Our approach outperforms existing methods by effectively utilizing pretrained segmentation models and cross-modal semantic alignment.
Index Terms— Audio-Visual Segmentation, SAM
@@ 1 Introduction
The Segment Anything Model (SAM) [5, 6] introduces an innovative approach to segmentation tasks by leveraging large-scale pretraining, similar to the way large language models use prompt-based learning [7]. The prompting idea has been applied to AVS. Mo and Tian [8] pioneered the idea of audio-prompted SAM by providing audio as a prompt to the SAM decoder. Liu et al. [9] advanced the audio-prompted SAM by introducing adapters into the image encoder of SAM. Audio-prompted SAM achieved better performance on the AVS task by leveraging the image segmentation foundation model, which mitigates the problem of the limited dataset. However, while SAM empowers the capacity of image understanding, the burden of learning audio-visual correspondence persists due to the scarcity of labeled data.
@@
Specifically, we introduce a novel feature, , which aligns audio and visual embeddings into a shared semantic space, leveraging the pre-trained cross-modal encoders CLIP [10] and CLAP [11]. By leveraging modality-specific projections, the model combines these embeddings through element-wise multiplication, capturing the intersection of audio and visual modalities. This design emphasizes shared semantics while reducing irrelevant information, producing a more robust and contextually enriched representation. Through this approach, our model effectively enhances audio-visual correspondence, enabling more accurate and reliable segmentation performance. Also, we report a vision bias problem in the single sound source dataset (S4). We found the vision bias from the dataset by performing better than previous state-of-the-art models without using audio information.
@@
We propose AV2T-SAM (Audio-Visual to Text SAM), the novel framework of utilizing a pre-trained text-prompted SAM by projecting audio features into the text embedding space. This strategy enables us to effectively harness the rich semantic information embedded in large-scale pre-trained cross-modal encoders, addressing the limitations posed by the scarcity of audio-visual segmentation masks.
We introduce a feature which captures the intersection of audio and visual modalities, emphasizing shared semantics and filtering out irrelevant information, resulting in a more contextually enriched and discriminative feature for segmentation tasks.
We demonstrate that the single sound source (S4) dataset contains a serious vision bias problem by surpassing previous state-of-the-art approaches without using any audio information.
@@ 2.2 Audio-Visual Segmentation
With the trends of multimodal learning, audio-visual tasks also got much attention. Aiming for better audio-visual correspondence, researchers worked on Sound Source Localization task [16, 17] that aims to localize the sound source portion from the frame. Due to the lack of a labeled dataset, researchers had to rely on Self-Supervised learning, which, while effective in leveraging unlabelled data, comes with limited applications to fine-grained tasks such as segmentation. Zhou et al. [3] introduce Audio-Visual Segmentation with a labeled AVSBench dataset. The AVSBench dataset enabled researchers to work on the Audio-Visual Segmentation task, which is more fine-grained than the localization task. With the new dataset, researchers started exploring audio-visual segmentation [3, 18]. Gao et al. [4] proposed an architecture that leverages the transformer in the AVS task. Liu et al. [19] introduced a new framework that separates the segmentation process and the verifying sounding object process.
@@ 3.1 Problem Formation
The input video of audio-visual segmentation (AVS) comprises of a series of visual frames and corresponding audio. The image frames are represented as V = , where () where denotes the -th visual frame with spatial dimensions , and is the total number of frames in the video. Each frame is paired with a corresponding 1-second audio waveform , where represents the number of audio samples in 1 second. The goal of the AVS task is to segment all sound source objects from each frame. The ground truth segmentation map is provided by binary masks where
@@ 3.2 Text-prompted Segmentation Module
The pre-trained text-prompted SAM model is used in our framework as a backbone to leverage its robust text-visual correspondence knowledge, which is from extensive text-image datasets [22, 23, 24]. While this module is interchangeable, we particularly used EVF-SAM [15]. The model employs the multimodal encoder, which is more efficient than adopting the Large Language model and performs better than the simple text encoder as the CLIP. While the original EVF-SAM model uses text as input, our method does not rely on explicit text. Instead, we map audio-visual features into the language space. To achieve this, we adapt their methods by removing the text encoder but keeping the multimodal encoder fixed, leveraging its strong adaptability to prompt SAM with text embeddings effectively.
3.3 Projection Module
This module consists of 2 Multi-layer Perceptron (MLP) projection layers that connect feature to the text embedding space required to prompt the pre-trained text-prompted segmentation model.
3.4 Semantically Aligned Features
@@
where and represent the transformed embeddings in the shared space. To capture the intersection of semantic information from both modalities, we calculate:
, where denotes the Hadamard (element-wise) product, and is the fused feature representation.
This operation ensures that the resulting feature emphasizes shared semantics between the two modalities while filtering out modality-specific noise. The design leverages the complementary nature of audio and visual information, providing a robust and contextually enriched representation that enhances downstream segmentation tasks. By projecting both modalities into a unified semantic space and fusing them through intersection, our approach effectively aligns and refines cross-modal information, addressing challenges in audio-visual correspondence learning.
@@ 3.5 Adapters
The mask decoder of SAM, which is two transformer layers, is too shallow to fuse the audio and visual information [9]. However, fine-tuning the image encoder of SAM requires considerable computing resources. To efficiently fuse audio and visual information while not entirely fine-tuning the image encoder, Liu et al. [9] suggested an audio adapter mechanism that trains only the adapter module which injects audio information to the frozen image encoder.
In this work, we employ the adapter module in the mask decoder of SAM with to effectively fuse prompt information to the image encoder. Specifically, for the -th layer of a transformer block, the -th adapter projects the prompt information to the output of the transformer block and repeats the prompt information along the spatial dimensions to match the dimensions of the image features, denoted as . The adapted features are then added to the output of the previous encoder layer, , to form the input to the -th layer, , as follows:
@@ 3.6 Learning Objectives
For training the model, we combined the Binary Cross-Entropy (BCE) loss and the Intersection over Union (IoU) loss, following the Liu et al. [9].
@@ 4.3 Implementation Details
We utilized EVF-SAM2 [15], which integrates SAM-2-L [6] as the SAM backbone and Beit-3-L [25] as the multimodal encoder for receiving a text prompt. For experiments using the SAM1 [5] as a backbone, we employed EVF-SAM1, which consists of SAM-H and Beit-3-L. The input image resolution was set to , and We trained the model for 40 epochs using the AdamW optimizer.
@@ 5.1 Main Results
| Category | Methods | Adapter | S4 (V1S) | MS3 (V1M) | ||
@@
We compare our model with state-of-the-art methods in Table 1. Our proposed model achieved new state-of-the-art performance on both subsets of ASVBench. On the S4 dataset, our model outperform the previous state-of-the-art, SAVE [21], with an improvement of 1.56 without introducing extra adapter layers into every image encoder block. Also, on the MS3 dataset, we exceed the previous state-of-the-art, ST-BAVA [20], without utilizing temporal information, which enabled ST-BAVA to outperform other approaches significantly.
@@
Figure 2 compares our segmentation results with those of existing models. In the left portion of the figure, we show one example of the results on the S4 dataset. TPAVI [3] captures only part of the guitar, suggesting weak visual object understanding. SAMA-AVS [9] leverages SAM to accurately segment the guitar but struggles to separate it from nearby objects. In contrast, our model fully segments the guitar and distinguishes it from its surroundings, made possible by strong audio-visual correspondence guided by the pretrained text-prompted SAM. A similar pattern emerges in the right portion of the figure, which depicts results on the MS3 dataset. Again, TPAVI exhibits partial segmentation of objects. While SAMA-AVS identifies the correct objects, it fails to separate the keyboard from a person. Our model, however, successfully segments the keyboard and isolates it from other objects. These findings underscore our model’s superior ability to detect and correlate audio with its corresponding visual region, thanks to the capabilities of the pre-trained text-prompted SAM.
@@ 5.2 Ablation on Adapters
5.3 Ablation on Feature
In this section, we compare the model’s performance with different prompt features. As denoted in a Table. 2, we can observe that perform best on both metrics over both datasets. This implies that feature, which captures information from both audio and visual modalities, contains richer information as the prompt than just the CLAP feature, which contains only audio information.
@@ 6 Conclusion
In this work, we have proposed a novel framework that leverages text modality embeddings and pre-trained text-prompted SAM on the audio-visual segmentation task. Also, we introduce a new feature that is aware of the intersection of audio and visual modalities for rich semantic information. Comprehensive experiment results have demonstrated that our approach performs comparable on the AVS task to state-of-the-arts models. Finally, we demonstrate the vision bias issue of the S4 dataset by surpassing previous state-of-the-art performance without using audio information.

### v2

@@ Abstract
The goal of Audio-Visual Segmentation (AVS) is to localize and segment the sounding source objects from video frames. Research on AVS suffers from data scarcity due to the high cost of fine-grained manual annotations. Recent works attempt to overcome the challenge of limited data by leveraging the vision foundation model, Segment Anything Model (SAM), prompting it with audio to enhance its ability to segment sounding source objects. While this approach alleviates the model’s burden on understanding visual modality by utilizing knowledge of pre-trained SAM, it does not address the fundamental challenge of learning audio-visual correspondence with limited data. To address this limitation, we propose AV2T-SAM, a novel framework that bridges audio features with the text embedding space of pre-trained text-prompted SAM. Our method leverages multimodal correspondence learned from rich text-image paired datasets to enhance audio-visual alignment. Furthermore, we introduce a novel feature, , which emphasizes shared semantics of audio and visual modalities while filtering irrelevant noise. Our approach outperforms existing methods on the AVSBench dataset by effectively utilizing pre-trained segmentation models and cross-modal semantic alignment. The source code is released at https://github.com/bok-bok/AV2T-SAM.
Index Terms— Audio-Visual Segmentation, Semantic Alignment, Text-Prompted Segment Anything Model
@@ 1 Introduction
The Segment Anything Model (SAM) [5, 6] introduces an innovative approach to segmentation tasks by leveraging large-scale pre-training, similar to the way large language models use prompt-based learning [7]. The prompting idea has been applied to AVS. Mo and Tian [8] pioneered the idea of audio-prompted SAM by providing audio as a prompt to the SAM decoder. Liu et al. [9] advanced the audio-prompted SAM by introducing adapters into the image encoder of SAM. Audio-prompted SAM achieved better performance on the AVS task by using the image segmentation foundation model, which mitigates the problem of the limited dataset. However, while SAM empowers the capacity of image understanding, the burden of learning audio-visual correspondence persists due to the scarcity of labeled data.
@@
Specifically, we introduce a novel feature, , which aligns audio and visual embeddings into a shared semantic space, leveraging the pre-trained cross-modal encoders CLIP [10] and CLAP [11]. By leveraging modality-specific projections, the model combines these embeddings through element-wise multiplication, capturing the intersection of audio and visual modalities. This design emphasizes shared semantics while reducing irrelevant information, producing a more robust and contextually enriched representation. Through this approach, our model effectively enhances audio-visual correspondence, enabling more accurate and reliable segmentation performance. We also identify a vision bias issue in the single sound source dataset (S4) of AVSBench, as our proposed system outperformed previous state-of-the-art models without using any audio information.
@@
We propose AV2T-SAM (Audio-Visual to Text SAM), a novel framework of utilizing a pre-trained text-prompted SAM by projecting audio and visual features into a text embedding space. This strategy enables us to effectively harness the rich semantic information embedded in large-scale pre-trained cross-modal encoders, addressing the limitations posed by the scarcity of audio-visual segmentation masks. Experimental results demonstrate the effectiveness of our method.
We introduce a feature that captures the intersection of audio and visual modalities, emphasizing shared semantics and filtering out irrelevant information, resulting in a more contextually enriched and discriminative feature for segmentation tasks.
We demonstrate a serious vision bias in the AVSBench S4 dataset by surpassing previous state-of-the-art approaches without using any audio information.
@@ 2.2 Audio-Visual Segmentation
With the trends of multimodal learning, audio-visual tasks also received much attention. Early researchers worked on the sound source localization task [16, 17] that aims to localize the sounding object in a frame. They often turned to self-supervised learning due to the lack of labels for localization. Though being effective to some extent, its application to fine-grained tasks like segmentation remains limited.
Zhou et al. [3] introduce Audio-Visual Segmentation (AVS) with a labeled AVSBench dataset containing fine-grained segmentation masks. With the new dataset, researchers began exploring audio-visual segmentation with supervised learning [3, 18]. Gao et al. [4] proposed a transformer-based architecture for the AVS task. Liu et al. [19] introduced a framework that separates the segmentation process and the process of verifying sounding object.
@@ 3.1 Problem Formation
The input video of audio-visual segmentation (AVS) comprises of a series of visual frames and corresponding audio. The image frames are represented as , where () where denotes the -th visual frame with spatial dimensions , and is the total number of frames in the video. Each frame is paired with a corresponding 1-second audio waveform , where represents the number of audio samples in 1 second. The goal of the AVS task is to segment all sound source objects from each frame. The ground truth segmentation map is provided by binary masks where
@@ 3.2 Text-prompted Segmentation Module
The pre-trained text-prompted SAM model is used in our framework as a backbone to leverage its robust text-visual correspondence knowledge, which is from extensive text-image datasets [22, 23, 24]. While this module is interchangeable, we particularly used EVF-SAM [15]. The model employs the multimodal encoder, which is more efficient than adopting the large language model and performs better than the simple text encoder, such as CLIP [10]. While the original EVF-SAM model uses text as input, our method does not rely on explicit text. Instead, we map audio-visual features into the language space. To achieve this, we adapt their methods by removing the text encoder but keeping the multimodal encoder fixed, leveraging its strong adaptability to prompt SAM with text embeddings effectively.
3.3 Semantically Aligned Feature Module
@@
To capture the intersection of semantic information from both modalities, we calculate:
where denotes the Hadamard (element-wise) product.
This operation ensures that the resulting feature emphasizes shared semantics between the two modalities while filtering out modality-specific noise.
Then, we pass the feature to the projection module, which consists of two MLP layers, to connect the features to the text embedding space required to prompt the pre-trained text-prompted segmentation model.
@@ 3.4 Adapters
Fine-tuning the image encoder of SAM requires considerable computing resources. To efficiently fuse audio and visual information while not entirely fine-tuning the image encoder, Liu et al. [9] suggested an audio adapter mechanism that trains only the adapter module which injects audio information to the frozen image encoder.
In this work, we employ the adapter module in the image encoder of SAM with to effectively fuse prompt information to the image encoder. Specifically, for the -th layer of a transformer block, the -th adapter projects the prompt information to the output of the transformer block and repeats the prompt information along the spatial dimensions to match the dimensions of the image features, denoted as . The adapted features are then added to the output of the previous encoder layer, , to form the input to the -th layer, , as follows:
@@ 3.5 Learning Objectives
For training the model, we combined the Binary Cross-Entropy (BCE) loss and the Intersection over Union (IoU) loss, following the Liu et al. [9],
@@ 4.3 Implementation Details
We utilize EVF-SAM2 [15], which integrates SAM-2-L [6] as the SAM backbone and Beit-3-L [25] as the multimodal encoder for receiving a text prompt. For experiments using the SAM1 [5] as a backbone, we employ EVF-SAM1, which consists of SAM-H and Beit-3-L. The input image resolution was set to , and We train the model for 40 epochs using the AdamW optimizer.
@@ 5.1 Main Results
| Category | Method | Adapter | S4 (V1S) | MS3 (V1M) | ||
@@
We compare our model with state-of-the-art methods in Table 1. Our proposed model achieves new state-of-the-art performance on both subsets of ASVBench. On the S4 dataset, our model outperform the previous state-of-the-art, SAVE [21], with an improvement of 1.56 without introducing extra adapter layers into every image encoder block. Also, on the MS3 dataset, we exceed the previous state-of-the-art, ST-BAVA [20], without utilizing their proposed temporal correspondence information.
@@
Figure 2 compares our segmentation results with those of existing models. In the left portion of the figure, we show one example of the results on the S4 dataset. TPAVI [3] captures only part of the guitar, suggesting weak visual object understanding. SAMA-AVS [9] leverages SAM to accurately segment the guitar but struggles to separate it from nearby objects. In contrast, our model fully segments the guitar and distinguishes it from its surroundings, made possible by strong audio-visual correspondence guided by the pretrained text-prompted SAM. A similar pattern emerges in the right portion of the figure, which depicts results on the MS3 dataset. Again, TPAVI exhibits partial segmentation of objects. While SAMA-AVS identifies the correct objects, it fails to separate the keyboard from a person. Our model, however, successfully segments the keyboard and isolates it from other objects.
These findings underscore our model’s superior ability to correlate audio with its visual region, thanks to our proposed semantically aligned feature and text-prompted SAM.
@@ 5.2 Ablation on Adapters
5.3 Ablation on Features
In this section, we compare the model’s performance with different prompt features. As denoted in a Table. 2, we can observe that performs best on both metrics over both datasets. This implies that feature, which captures information from both audio and visual modalities, contains richer information as the prompt than just the CLAP feature, which contains only audio information.
@@ 6 Conclusion
In this work, we proposed a novel framework that leverages text modality embeddings and pre-trained text-prompted SAM on the audio-visual segmentation task. Also, we introduced a new feature that is aware of the intersection of audio and visual modalities for rich semantic information. Comprehensive experiment results demonstrated that our approach performs comparable on the AVS task to state-of-the-arts models. Finally, we demonstrated the vision bias issue of the S4 dataset by surpassing previous state-of-the-art performance without using audio information.
7 Acknowledgments
You Zhang would like to thank the synergistic activities provided by the NRT program on AR/VR funded by NSF grant #1922591. The GPU resources are supported by NAIRR Pilot Project #240152, ACCESS #ELE240019, and NCSA Delta.